OpenStudy (anonymous):

Solve: x + 3 = −x + 7

OpenStudy (ghazi):

2x= 7-4

OpenStudy (ghazi):

x=3/2

OpenStudy (anonymous):

x+3=-X+7 2x+3=7 2x=4 x=2